# Automatic infant 2D pose estimation from videos: comparing seven deep neural network methods

### v2

@@ Results
The complete table with the percentage of redundant detections is shown in Tab. ST. 7 and ST. 8 for real and synthetic infants respectively.
For real infants, it is difficult to estimate the difficulty of each sequence exactly. We observe that video input has a tendency to produce more redundant detections.
For synthetic infants, we observe that Detectron2 and HRNet RD seem to produce more redundant detections when the sequence is difficult (IDs 10-12), while HRNet BU seems to produce more redundant detections when the sequence is easy (IDs 1-4). We observe a tendency to produce more redundant detections with video input than for image input.
Real Video ID AlphaPose DeepLabCut Detectron 2 MediaPipe HRNet BU HRNet TD OpenPose ViTPose
Images AA_8w 0 N/A 11.5 0 1.4 1.1 0 N/A
AA_17w 0.6 N/A 8.0 0 1.1 40.5 0 N/A
TH_8w_st1 0 N/A 9.9 0 0.5 3.3 0 N/A
TH_8w_st2 0 N/A 34.6 0 5.3 22.1 0 N/A
TH_8w_st3 1.4 N/A 34.3 0 2.6 12.1 0 N/A
TH_15w 14.4 N/A 86.3 0 11.8 103.5 0 N/A
TH_19w 0.4 N/A 28.9 0 7.8 53.2 0 N/A
TH_25w 0.4 N/A 7.1 0 5.1 54.6 0 N/A
Videos AA_8w 0 0 92.0 0 9 1.0 0 1.4
AA_11w 0.3 0 199.8 0 0.3 44.9 0 0.1
AA_13w 4.7 0 290.9 0 4.9 66.7 0 4.0
AA_17w 1.2 0 170.1 0 1.4 47.7 0 6.4
AA_19w 6.7 0 308.6 0 10.0 26.4 0 4.0
TH_8w_st1 0.1 0 197.0 0 0.9 1.9 0 2.9
TH_8w_st2 0 0 188.7 0 6.5 20.6 0 38.6
TH_8w_st3 0.9 0 249.2 0 2.9 12.7 0 26.6
TH_10w_st1 0.1 0 197.0 0 0.1 7.7 0 22.5
TH_10w_st2 1.2 0 388.0 0 3.3 130.0 0 82.3
TH_12w 0.1 0 164.4 0 0.3 5.4 0 5.4
TH_15w 15.2 0 522.8 0 13.9 98.8 0 28.8
TH_17w 2.4 0 198.9 0 9.4 101.2 0 40.3
TH_19w 0.7 0 257.6 0 8.2 51.5 0 2.7
TH_23w 51.9 0 198.0 0 2.3 91.0 0 43.9
TH_25w 0.5 0 211.3 0 4.9 43.1 0 36.6
Table 7: Percentage of redundant detections for each method and input type on real infants videos.
@@
Synth. Video ID AlphaPose DeepLabCut Detectron 2 MediaPipe HRNet BU HRNet TD OpenPose ViTPose
Images Syn. 1 0.3 N/A 5.5 0 56.8 0.6 0 N/A
Syn. 2 0.5 N/A 1.3 0 0 0.6 0 N/A
Syn. 3 0 N/A 0 0 0 0.2 0 N/A
Syn. 4 0 N/A 0.1 0 20.7 3.6 0 N/A
Syn. 5 0 N/A 0 0 3.6 1.6 0 N/A
Syn. 6 0 N/A 0 0 0 7.1 0 N/A
Syn. 7 0 N/A 2.0 0 0.1 0.3 0 N/A
Syn. 8 0 N/A 0.5 0 1.9 0.5 0 N/A
Syn. 9 0 N/A 29.0 0 6.2 95.4 0 N/A
Syn. 10 0.1 N/A 240.2 0 1.9 101.6 0 N/A
Syn. 11 0 N/A 102.2 0 1.0 0.4 0 N/A
Syn. 12 0 N/A 31.7 0 7.6 106.8 0 N/A
Videos Syn. 1 0.5 0 20.5 0 61.8 1.5 0 2.6
Syn. 2 0 0 38.1 0 0 2.7 0 0
Syn. 3 0 0 0 0 0.1 0.1 0 0
Syn. 4 1 0 92.2 0 31.3 15.1 0 2.4
Syn. 5 0 0 13.6 0 9.1 3.7 0 3.6
Syn. 6 0 0 42.2 0 0.9 24.0 0 9.2
Syn. 7 0 0 98.7 0 1.3 0.9 0 0.1
Syn. 8 0 0 70.7 0 2.9 2.2 0 0
Syn. 9 0.1 0 97.2 0 23.7 93.1 0 50.4
Syn. 10 0.4 0 378.1 0 3.3 30.6 0 2.5
Syn. 11 0 0 137.1 0 1.0 0.8 0 74.5
Syn. 12 0.1 0 121.6 0 5.8 110.5 0 30.5
Table 8: Percentage of redundant detections for each method and input type on synthetic infants.
@@
In a best-case scenario, 3, pose estimation methods car reach high levels of performance on infants in supine position, with low variability, with ViTPose reaching an Average Precision up to 91.8, and average errors around 3.9$\pm$2.0of the Neck-MidHip segment (which corresponds roughly to the infant’s torso length).
With the detailed OKS values for each video (Tables ST. 4 and 5), we could get further insight with regards to the complexity estimation made by Hesse et al. on their MINI-RGBD dataset based on the actual performance of the methods’ estimations. Such table can also help to identify for each specific method which kind of sequences they seem to struggle with to identify the possible causes of keypoints misplacement, so that more of such examples can be included in future methods’ training or fine-tuning. For example, MediaPipe did not manage to identify a single image among the sequence for synthetic infant 2, despite it being deemed "easy", possibly due to its unique background.
From Table 6 of the main manuscripts and Tables of Supplementary Materials ST 2 and ST 3, it seems that the better the estimations and the easier the videos, the lower the correlation between scores and OKS values. This could be explained by a lower range of high values from OKS, while the scores might be less reliable and might not reflect a similar reduction in their variability and range of values.

### v3

@@ Results
Missing Data
The percentages of missing data are shown in Tab. 5.
For real infants, the methods that never miss detections or keypoints are DeepLabCut and Detectron2, while HRNet BU is very close behind, with only a few misses that are lost when rounded at less than 2 decimal points. ViTPose and HRNet TD miss an amount of data that could be considered acceptable, below 0.5%. On the other hand, OpenPose, MediaPipe, and AlphaPose display high amounts of missing data. We observe a tendency to miss smaller amounts of data with video input compared to image input. OpenPose reaches more than 50%, the highest number of missing data overall, on the Lap dataset.
For synthetic infants, the methods missing data are OpenPose, MediaPipe, and AlphaPose, though to a lesser degree than for real infants. Contrary to the real infants, we observe a tendency to miss smaller amounts of data with image input compared to video input.
Real Images 7.4 N/A 0 9.3 0 0.1 10.1 N/A
Supine Videos 6.9 0 0 9.5 0 2.9 8.3 0.3
[…]
Synth. Images 5.7 N/A 0 8.5 0 0 3.6 N/A
MINI-RGBD Videos 9.9 0 0 8.5 0 0 3.7 0
@@
The percentages of redundant detections for each method are shown in Tab. 6. The details for each individual video can be found in the Supplementary Tables ST. 11 and 12. DeepLabCut, MediaPipe and OpenPose cannot have redundant detections by design, though for the case of the ”Lap” dataset, they could have minus detections. For the other methods, on the ”supine” real infants dataset, the methods with the least amount of redundant data are AlphaPose and HRNet BU. Except for HRNet TD, we observe a tendency, especially for Detectron2, to produce more redundant detections with video input. On the ”Lap” dataset, HRNet BU, Detectron2, and HRNet TD display higher rates of redundant detections, while on the contrary, MediaPipe, ViTPose, and AlphaPose seem to not always detect both individuals and only detect one of them (often the adult) for high percentages of the images. OpenPose is the method that overall has the least difference in detections compared to the expected amount. On SyRIP, all methods except MediaPipe and OpenPose display high amounts of redundant detections.
For synthetic infants, AlphaPose has the least amount of redundant detections, outside of the methods that could not have any.
Overall, Detectron2 and HRNet TD overall provide disproportionate amounts of redundant detections, especially on real infants, though HRNet BU has the most redundant detections in the ”Lap” context, where, on the contrary, MediaPipe struggles the most at providing at least 2 detections (one of each individual on the image).
In the ”supine” and MINI-RGBD datasets, it is possible to look at the details per video (see Supplementary Materials Tables ST 11 and 12.) In doing so, we see that the distribution of redundant detections is not uniform. Detectron2, HRNet TD and ViTPose show higher redundant detections for synthetic infants that are considered more difficult in the MINI-RGBD dataset, while HRNet BU shows higher redundant detections for the easier synthetic infants.
Real Images 2.5 N/A 32.4 0 6.2 51.4 0 N/A
Supine Videos 2.7 0 277.1 0 6.7 48.8 0 23.1
[…]
Synth. Images 0 N/A 34.4 0 8.3 26.6 0 N/A
MINI-RGBD Videos 0.1 0 92.5 0 11.8 23.8 0 14.7
Correlations between scores and OKS
The different methods provide internal estimates of the quality of their detections in each frame (score for the whole keypoint set; confidence for each keypoint). This can be useful for downstream tasks—detections with a low score can be ignored, for example. We investigated the correlation between scores (method estimates) and actual accuracy (OKS values). Table  7 reports the Spearman Rank Correlation Coefficients.
For real infants, in the ”supine” context, the method with the highest Correlation Coefficients is HRNet BU for both input types. With image input, it is closely followed by OpenPose. The correlations are in the correct direction for all methods. It is interesting to note that on the ”Lap” dataset, some methods maintain correlation scores in line with the other datasets, while others have lower correlations (AlphaPose, OpenPose, and ViTPose). In contrast, on the SyRIP dataset, most methods display higher correlations, especially AlphaPose. Still, HRNet BU is the method with the highest correlations in all the real infants datasets. For synthetic infants, the methods with the highest Correlation Coefficients is OpenPose for image inputs. For video inputs, HRNet BU has the highest correlations, closely followed by OpenPose.
Overall, the correlations between the scores and the real accuracy are moderate, as the highest reach values of 0.68, while most of the correlations are low, between 0.2 and 0.5, and should therefore be used with caution.
For completeness, the scatterplots of the score and OKS values for each method on each dataset are in the Supplementary Materials, Figs. SF. 3-6.
Real Images 0.28 N/A 0.19 N/A 0.68 0.47 0.63 N/A
Supine Videos 0.31 0.01 ($p=0.79$) 0.19 N/A 0.67 0.49 0.28 0.52
[…]
Synth. Images 0.41 N/A 0.51 N/A 0.29 0.43 0.58 N/A
MINI-RGBD Videos 0.39 0.46 0.53 N/A 0.66 0.51 0.61 0.48
Processing speed
Table 8 shows the processing speeds for each method in the synthetic infant dataset. DeepLabCut and AlphaPose are the fastest methods by a large margin, running at close to 30 fps, while MediaPipe and OpenPose are close to running at 15 fps.
@@
Synth. AlphaPose DeepLabCut Detectron 2 MediaPipe HRnet BU HRnet TD OpenPose ViTPose
Images 27.5 N/A 7.1 14.1 2.4 6.5 14.0 N/A
Videos 27.0 28.8 5.9 15.4 2.3 7.1 13.0 4.8
Conclusion
With extremely rapid progress in human pose estimation methods from images and videos, this technology lends itself to deployment in infant motion analysis, providing a tool that can be truly applied “in the wild” with inputs, say, from a cell phone camera and without additional requirements or restrictions. If the pose estimation accuracy is satisfactory, this will have far-reaching implications for our understanding of normal motor development as well as the early diagnosis of infant developmental disorders.
This article provided an empirical comparison of seven state-of-the-art human pose estimation methods—trained on images of adults during various activities, typically in upright positions—on several datasets of infant images, including infants under 7 months of age in supine position, infants between 17 weeks to 18 months in close proximity with a parent, and infants below 1 year old in a variety of contexts and postures. Our results are summarized below.
First, we conclude that state-of-the-art human pose estimation methods work well to estimate infant poses without the need for additional training or fine-tuning for infants in ”supine” position. An overview of the estimation accuracy for individual keypoints on the body is provided in Fig. 2, and in Supplementary Materials Tables ST 6-9. Overall accuracy estimations are available in Tables  2,  3 using standard performance metrics (OKS, AP/AR) and in Table 4 using an infant-size centred metric. ViTPose has the best accuracy [49], followed by HRNet [47, 48] (top-down variant). The other networks tested (HRNet bottom-up, AlphaPose [31], Detectron2 [44], and OpenPose [28]) have higher errors. The DeepLabCut environment [51], with human pose estimation using DeeperCut [43], as well as MediaPipe [45] with BlazePose [46] does not provide competitive results at all. It is coherent that ViTPose, being based on one of the newest state-of-the-art neural network architecture, Transformers, which introduce an ”attention” mechanism, as well as being trained on more datasets that the other methods, ended up with the best performance. These results are in line with [42] who used the Percentage of Correct Keypoints (PCK), an accuracy-based evaluation that can be compared to our Neck-MidHip errors, as they also base it on the infant’s approximate torso length. The performance rank order of the methods we commonly compare are the same, with the identical conclusion that ViTPose performs the best. The error levels also converge towards similar levels: they find that 84.6% of the ViTPose estimations are equal or below 10% of the infant torso length, while we find that 68% of the ViTPose estimations are within 3.3% to 8.7% of the infant torso length. However, compared to [42], who only used a single accuracy-based evaluation, PCK, we complemented our results using several accuracy-based metrics, including the standard AP and AR, with other analyses such as the amount of missing data or of produced redundant detection, the relation between scores and confidence values with accuracy-based performance, as well as the processing speed of each method. We also included some methods that we observed to be often used by infant development researchers (AlphaPose and DeepLabCut in particular).
On the more complex ”Lap” and SyRIP datasets, ViTPose and either OpenPose or HRNet depending on the metric and dataset have the highest performance. However, for all methods, the drop in performance is drastic, for example losing between 32 to 65 points in Average Precision and no method reaching errors lower than an average of 12% of the infants’ torso length.
Second, we complemented accuracy-based performance analysis with additional criteria that are important for the practical applications of the methods. The accuracy (OKS, Neck-midhip errors) cannot be computed if keypoints are not detected and does not penalize the case where multiple people are erroneously detected (though AP partially consider these cases). Moreover, both missing and redundant detections negatively impact any downstream motion analysis. Detectron2 and HRNet (top-down) are in particular susceptible to redundant detections. OpenPose, MediaPipe and AlphaPose have high rates of missing data (missed detections of the whole infant or of individual keypoints). In the case of infant-parent (dyadic) interaction settings, such as in our ”Lap” dataset, disentangling individual detections is not trivial, even when using dual RGB-D camera setup in a laboratory setting [59]. One reason for this is that the methods are biased towards adults instead of infants, and will sometimes only detect the parent. This is particularly noticeable when looking at methods that are unable to provide as many detections as expected in this context; see Table  6.
It is also important to consider both accuracy-based metrics (AP/AR, Neck-MidHip, OSK) together with higher level metrics (missing data, redundant detections): for example, OpenPose, MediaPipe and AlphaPose have high amounts of missing data. Missing keypoints, in particular, which happen for bottom-up methods either because the methods cannot detect them in the image, or because their tentative placement was too far from the model’s priors and were ignored, can not be evaluated by accuracy-based errors. In contrast, top-down methods would have a tendency to provide all keypoints even when misplaced, giving a lower confidence or score value to flag this. While confidence thresholds can be arbitrarily chosen to cull and remove those keypoints, which some methods do internally, we decided to not set such an arbitrary threshold and to not remove them in our results for the methods that did not, making these keypoints impact the accuracy-based metrics instead of showing as missing data. The most striking example might be OpenPose, a Bottom-Up method, which misses up to 54% of the data, mainly as missing keypoints, in the ”Lap” dataset, yet manages to reach the second highest AP values and the second lowest average Neck-MidHip errors for this dataset.
Third, we evaluated the computation time of all the methods on the same computer. On this machine, which can be considered a powerful workstation in 2024, most of the methods could process less than 15 frames per second. AlphaPose was an exception; with approximately 27 fps, it could be deployed if a real-time pose estimation is necessary. On the other hand, any application that does not need real-time processing would not care about this, and rather focus on the most accurate method, ViTPose, even if it is much slower.
Discussion and Future Work
Researchers in behavioral science are not experts in machine learning and computer vision and cannot keep up with the rapid progress in human pose estimation methods. Hence, they choose their tools based on methods that have been used in the past (in the context of infants, OpenPose was used in [29, 33, 30]; AlphaPose in [20]). The DeepLabCut environment [51] is the frequent tool of choice in behavioral science. Here we show that for the estimation of the pose of infants in supine position, DeepLabCut with human pose estimation using DeeperCut [43] currently does not provide satisfactory performance at all. OpenPose and AlphaPose may be employed but fall behind ViTPose and HRNet (top-down) in accuracy, as well as missed data. We recommend the community to use the currently best-performing tools and profit from the Docker containers we released.
We would like to emphasize that the human pose estimation tools are unfortunately not standard computer programs that can be easily deployed and used “as is”. Their installation may be more or less involved. Moreover, performance is affected by preprocessing steps and settings. For example, cropping images so that only the infant is in the image helps to improve performances. Images with multiple persons in the scene (e.g., caregiver or experimenter), especially in close proximity, are currently challenging, and body keypoints of adults and the infant can even get mixed up. Also, as the pose estimation networks were predominantly trained on people in upright positions, the orientation of the infant image importantly affects the performance—the camera should be positioned or every frame rotated afterwards so that the infant is positioned approximately with the head at the top. Some pose estimation networks may have additional settings (e.g., OpenPose and MediaPipe allow to set the maximum number of persons to detect in the image).
Pose estimation methods can take either videos or individual images as input (or both). Although this may seem equivalent at first sight—a video is a sequence of consecutive images—, we realized that we sometimes obtained quite different results for these two input modes. The differences might partially come from how each method cuts the videos into frames, which can lead to slight offsets compared to the ground-truth sequences we used for evaluation. However, this alone cannot explain the larger differences in performance for OpenPose or the different in the number of redundant detections for Detectron 2. Some of the pose estimation methods have divergent processing paths for image and video processing. In particular, tracking, a step that exploits the temporal consistency of estimations over the sequence of detections, when implemented, is often enabled only for video input and not for images. With the exception of MediaPipe, for which we enabled it for both input modes in our ”supine” and MINI-RGBD datasets thanks to available parameters, none of the methods we tested had a clear tracking option in its parameters. Tracking can be currently achieved in postprocessing, as not all methods contain such a step, but in the future it is likely to find its way into the pose estimation methods themselves.
The top-down methods, which identify areas (bounding boxes) where persons are likely to be in the images first, and then look for keypoints on the body within these areas, require a (bounding-box) detector that provides the coordinates of this bounding box, in theory encompassing all the visible body parts of detected individuals. They are the main source of redundant and missed detections for these methods. Detectors are separate from the methods but often embedded within them, and some work might be needed to compare their performance in detecting infants. For example, we hypothesize that the poor performance of MediaPipe is influenced by its detector, which was made to detect (adult) faces instead of bodies and assumes that the body is directly below, as a compromise to increase processing speed. We can also observe that on our ”Lap” dataset, it is the bottom-up methods, HRNet BU and OpenPose that have the lowest Neck-MidHip errors, while all top-down methods show much higher errors. Another example is by looking at ViTPose’s reported results on the OCHuman dataset: a dataset comprised of multi-body scenes with adults. The AP for ViTPose-b is reported to be 87.3 when using the ground truth bounding boxes ([49] Table 11), while another work comparing the performance of several detectors on this dataset showed that the AP of ViTPose-b using the RTMDet detector drops to 45.0; and even the best detector only reached 48.3 AP on this dataset ([60], Table 1.
In this work, the “vanilla” versions of pose estimation methods, the models trained on datasets featuring mostly adults, were directly used to process videos with infants in supine position. Contrary to expectations, the drop in performance was not dramatic. In fact, comparing the average precision and recall (AP, AR) results with those reported for these methods on the COCO validation or the test datasets, or on the “body” part of the Halpe and COCO-WholeBody datasets [49, 31, 47, 48], it seems that VitPose performs better on real infants in supine position than reported on adults by around 9 and 7 points respectively. OpenPose also performs better than reported by around 10 points in both AP and AR when processing images, but worse when processing videos. For AlphaPose and HRNet BU, the results are roughly similar. For HRNet TD, its AP is at least 10 points lower, most likely due to higher amounts of high-scoring redundant detections, although its AR is around 8 points higher.
Such results could be explained by the dataset containing infants only in supine position with a view from above, which, as long as the infants do not perform complex leg movements, is rather simple and could be considered similar enough to an adult standing in front of a camera: the situation that these methods encounter most commonly during their training. This is supported by looking at the average OKS values from individual videos (Supp. Materials ST. 6), where the performance of all methods decreases noticeably and is more variable for “hard” synthetic infants (IDs 10-12) compared to their performance on “easy” infants (IDs 1-4). Furthermore, using more complex datasets, such as SyRIP or our ”Lap” dataset shows large drops of performance at all levels. The ”Lap” condition, which would fall under a multi-body detection scenario, is challenging even when only adults are involved, partially due to the poor performance of detectors for top-down methods [60]. In our results, ViTPose-h performs 15 AP points lower in our ”Lap” dataset compared to ViTPose-b on the OCHuman dataset, which might highlight a bias toward detecting adults rather than infants when both are present in a scene together due to the training datasets only containing adults. It also scores about 40 points lower in AP on the SyRIP dataset compared to the common benchmark dataset for adults, COCO.
There are several directions for future work. First, although small infants (under seven months) differ significantly in their body proportions from adults and are thus strongly “out of distribution” data for the models in terms of shape estimation, pose estimation may be relatively easier because infants are in supine position all the time and their motor repertoire is constrained. Older infants and children in prone, standing, and other postures which are partially covered in the SyRIP dataset should be considered as a future challenge for pose estimation, and be included in the training of future methods considering the difficulty for current methods to handle such postures. The situation is similar for scenes with multiple people, like infants on a parent’s lap. Second, in addition to 2D pose estimation in the form of keypoints, there are methods that estimate 3D pose and shape (a complete body mesh) from an image or video (e.g., Smplify-x [58], with SMIL for an infant shape [61], or 4D Humans currently with adult shape only [62]). Third, while pose estimation is a key prerequisite for additional analyzes, the quality of motion extracted from the sequences of estimated keypoints should be studied in detail (see also [63]).
Unlike for adults, datasets of infant video recordings with additional reference (e.g., from motion capture, or manually-annotated) are scarce and often kept private. We are open to collaborate with research groups that are willing to share such datasets. We would also be happy to share our experience on how to best record infants such that the pose estimation result is optimal.
[…]
Supplementary Information This work has supplementary materials that include additional details about the pose estimation methods and datasets as well as additional results and analysis.
[…]
Acknowledgements We thank Miroslav Purkrábek and Pramod Murthy for their advice, and Lukáš Rustler, Valentin Marcel, Jason Khoury, Vojtěch Ježek, and Vojtěch Volprecht for their assistance.
Declarations
Funding
This work was supported by the Czech Science Foundation (GA CR), project no. 20-24186X and project no. 25-18113S. F.G. was additionally supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS24/096/OHK3/2T/13. S.T.P was additionally supported by the project Mobility ČVUT MSCA-F-CZ-I CZ.02.01.01/00/22_010/0003405.
Conflicts of interest/Competing interests
The authors declare no competing interests.
Ethics approval and consent to participate and for publication
The current study used video recordings of two full-term healthy infants (1f, 1m), observed between 8 weeks and and 25 weeks of age in supine position, and two full-term healthy infants (2f) observed at 4 months and 1 year and 3 months of age sitting on the lap of a parent. Informed consent was obtained from the participants’ legal guardians, including consent for publication of images in an online open-access publication. The study was approved by the Committee for Research Ethics at the Czech Technical University in Prague under reference number 00000-07/21/51903/EKCVUT and was carried out in accordance with all relevant guidelines and regulations. We additionally included video recordings of four full-term infants (2f, 2m), observed between 6 and 18 months of age sitting or standing in front of their parents. Informed consent was obtained from the participants’ legal guardians. The study was approved by the Pavia Ethics Committee (p–20190004177).
Availability of data, code, and materials
Two of the datasets used (MINI-RGBD dataset [54] and SyRIP dataset [55]) are publicly available in their respective public repositories. Video recordings and images from our infant datasets (“supine” and “lap”) cannot be publicly shared due to data protection and sharing rules.
We publicly share the Docker images and Dockerfiles of the pose estimation methods that we created and used, all the program scripts used to compute the analyses, as well as the anonymized sections of our datasets (estimated keypoints) and the analysis output files at the following locations: https://osf.io/x465b/ and https://hub.docker.com/u/humanoidsctu. DeepLabCut’s official installation instructions include creating its own conda environment, while MediaPipe is already its own easy-to-install Python library, and as such did not need to be further containerized with Docker, although we provide Python scripts and instructions for utility. Due to a technical error during upload, the OpenPose container could not be uploaded to DockerHub at the time of writing. We do not share the direct outputs of the methods due to the sheer amount of files, but their content is available in the provided resources in a compact and easy-to-use format.
Author contribution
The work was conceived by F.G. and M.H. The data was collected by M.H., L.N. and F.G. The data was processed by F.G, M.M. and L.N. The data was analyzed by F.G. The manuscript was prepared by F.G. and revised and reviewed by M.H. and S.T.P. Funding was acquired by M.H.
@@ Conclusion
In a best-case scenario (see Tab. S 4), pose estimation methods can reach high levels of performance on infants in supine position, with low variability, with ViTPose reaching an Average Precision up to 91.8, and average errors around 3.9$\pm$2.0of the Neck-MidHip segment (which corresponds roughly to the infant’s torso length).
With the detailed OKS values for each video (Tables ST. 5 and 6), we could get further insight with regards to the complexity estimation made by Hesse et al. on their MINI-RGBD dataset based on the actual performance of the methods’ estimations. Such a table can also help to identify for each specific method which kind of sequences they seem to struggle with to identify the possible causes of keypoint misplacement, so that more of such examples can be included in future methods’ training or fine-tuning. For example, MediaPipe did not manage to identify a single image among the sequence for synthetic infant 2, despite it being deemed ”easy”, possibly due to its unique background.
From Table 7 of the main manuscripts and Tables of Supplementary Materials ST 3 and ST 4, it seems that the better the estimations and the easier the videos, the lower the correlation between scores and OKS values. This could be explained by a lower range of high values from OKS, while the scores might be less reliable and might not reflect a similar reduction in their variability and range of values.